Larson, Inc., manufactures backpacks. Last year, it sold 77,500 of its basic model for $15 per unit. The company estimates that this volume represents a 25 percent share of the current market. The market is expected to increase by 10 percent next year. Marketing specialists have determined that as a result of new competition, the company’s market share will fall to 20 percent (of this larger market). Due to changes in prices, the new price for the backpacks will be $12 per unit. This new price is expected to be in line with the competition and have no effect on the volume estimates.

Required:

Estimate Larson’s sales revenues from this model of backpack for the coming year.

Sales revenue -
Answer #1

Sales of current year = 35%

Increase in sales = 15%

Companies market share falls to 20%

Sales revenue = 77,500 X 110/25 X 20/100 X $12 = $818,400
Answer #2

step 1) 77,500/25%= 310,000

step 2) 310,000*10%= 31,000

step 3) 310,000+31,000= 341,000

step 4) 341,000*20%= 68,200

step 5) 68,200*12= 818,400